OpenStudy (anonymous):

The dimensions of a smaller rectangle are 3 ft by 9 ft. The dimensions of a larger rectangle are 5 ft by 15 ft. Find the ratio of the area of the smaller rectangle to the area of the larger rectangle

OpenStudy (anonymous):

Area #1 = 3*9 =27 area 2= 5*15=75. Ratio 27:75 reduces to 9:25